3.60 x 102 V 40.0 V - 80.0 - 80.0 V 2.00 x 10 NŠ 80.003 20.093 70.0 (a) the current (in A) in each resistor 1200 magnitude diType or paste question here
Answer #1

20v low- (2 & 64 = 70 R, is too 1, to loop 2 VAR, -y figh- DA Tì, 1 toi - - - e liq-5,1-0 oth (tigtin) - 360-40 + 8012 Eve :.14: +86 12 = 242 2.71, +42122136 0x = + (1 xs, 1412 & 43x5z = 7+ 13645 12 (228) = 677 :.= $(21, 4) CIA = { [40-21,- tois) at= R - 2000 Po = Voltage ) (current possing through it) - yox iz : 120 W 6: 360*Ciptle tig) = 36078 - 2880 & Po = sox - 320w